# Learning a Neural Solver for Parametric PDE to Enhance Physics-Informed Methods

### v2

@@ Abstract
Physics-informed deep learning often faces optimization challenges due to the complexity of solving partial differential equations (PDEs), which involve exploring large solution spaces, require numerous iterations, and can lead to unstable training. These challenges arise particularly from the ill-conditioning of the optimization problem, caused by the differential terms in the loss function. To address these issues, we propose learning a solver, i.e., solving PDEs using a physics-informed iterative algorithm trained on data. Our method learns to condition a gradient descent algorithm that automatically adapts to each PDE instance, significantly accelerating and stabilizing the optimization process and enabling faster convergence of physics-aware models. Furthermore, while traditional physics-informed methods solve for a single PDE instance, our approach addresses parametric PDEs. Specifically, our method integrates the physical loss gradient with the PDE parameters to solve over a distribution of PDE parameters, including coefficients, initial conditions, or boundary conditions. We demonstrate the effectiveness of our method through empirical experiments on multiple datasets, comparing training and test-time optimization performance  [^1] Code will be released soon at https://github.com/2ailesB/neural-solver.
@@ 1 Introduction
Partial Differential Equations (PDEs) are ubiquitous as mathematical models of dynamical phenomena in science and engineering. Solving PDEs is of crucial interest to researchers and engineers, leading to a huge literature on this subject (Evans, 2010; Salsa, 2015). Traditional approaches to solving PDEs such as finite difference, finite element analysis or spectral methods (Zienkiewicz et al., 2005; LeVeque, 2007) often come with stability and convergence guarantees, but suffer from a high computational cost. Improving numerical PDE solvers through faster and more accurate algorithms remains an active research topic (Zienkiewicz et al., 2005).
@@
PDE solvers usually rely on discretization and/or linearization of the problem through various techniques, to simplify the computations. Iterative methods such as Jacobi, Gauss-Seidel, Conjugate Gradient, and Krylov subspace methods can then be used to solve the resulting systems. Unfortunately, many PDEs have an ill-conditioned nature and these iterative processes can demand extensive computational resources. Preconditioning techniques are often essential to mitigate this, though they require precise customization to the specific PDE problem, making the development of effective solvers a significant research endeavor in itself. Yet, the computational demands, time, and expertise required to develop these algorithms sometimes make them impractical or sub-optimal for specific classes of problems. Instead of relying on hand-designed algorithms, researchers have investigated, as an alternative, the use of machine learning for training iterative PDE solvers Hsieh et al. (2019); Li et al. (2023a); Rudikov et al. (2024); Kopaničáková et al. (2023). These approaches usually parallel the classical numerical methods by solving a linear system resulting from the discretization of a PDE, for example, using finite differences or finite elements. A preconditioner is learned from data by optimizing a residual loss computed w.r.t. a ground truth solution obtained with a PDE solver. This preconditioner is used on top of a baseline iterative solver and aims at accelerating its convergence. Examples of baselines solvers are the conjugate gradient (Li et al., 2023a; Rudikov et al., 2024) or the Jacobi method Hsieh et al. (2019).
@@
Another recent research direction investigates the use of neural networks for building surrogate models in order to accelerate the computations traditionally handled by numerical techniques. These methods fall into two main categories: supervised and unsupervised. The supervised methodology consists of first solving the PDE using numerical methods to generate input and target data and then regressing the solution using neural networks in the hope that this surrogate could solve new instances of the PDE. Many models, such as Neural Operators, lie within this class (Li et al., 2020; Raonić et al., 2023; Bartolucci et al., 2023) and focus on learning the solution operator directly through a single neural network pass. Unsupervised approaches, involve considering a neural network as a solution of the PDE. The neural network parameters are found by minimizing the PDE residual with gradient descent. Methods such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), or DeepRitz (E & Yu, 2018) fall under this category. This family of methods is attractive as it does not rely on any form of data, but only on information from the PDE residual. However, they exhibit severe difficulties for training (Krishnapriyan et al., 2021; Ryck et al., 2023), often requiring many optimization steps and sophisticated training schemes (Krishnapriyan et al., 2021; Rathore et al., 2024). The ill-conditioned nature of PDEs residual loss appears again in this context, making standard optimizers such as Adam inappropriate. A more detailed review of the existing literature is described in appendix A.
@@
In this work, we consider having access to the PDE as in unsupervised approaches and also to some data for training our neural solver. Our objective is to solve the optimization issues mentioned above by learning an iterative algorithm that solves the PDE from its residual defined as in the PINNs framework (see fig. 1). This neural solver is trained from data, either simulations or observations. Different from the classical ML training problem which aims at learning the parameters of a statistical model from examples, the problem we handle is learning to learn Andrychowicz et al. (2016), i.e. learning an iterative algorithm that will allow us to solve a learning problem. When vanilla PINNs handle a single PDE instance, requiring retraining for each new instance, we consider the more complex setting of solving parametric PDEs, the parameters may include boundary or initial conditions, forcing terms, and PDE coefficients. Each specific instance of the PDE, sampled from the PDE parameter distribution, will then be considered as a training example. The objective is then to learn a solver from a sample of the parametric PDE distribution in order to accelerate inference on new instances of the equation. With respect to unsupervised approaches, our model implements an improved optimization algorithm, tailored to the parametric PDE problem at hand instead of using a hand-defined method such as stochastic gradient descent (SGD) or Adam. As will be seen in the experimental section, this idea will prove extremely effective for the ill-posed problem of optimizing PINNs objective, allowing convergence in a few steps. In the proposed methodology, the neural solver will make use of gradient information computed by a baseline gradient method in order to accelerate its convergence. In our instantiation, we will use SGD as our baseline algorithm, but the method could be easily extended to other baselines. Our model deviates from the traditional preconditioning methods by directly optimizing the non-linear PDE residual loss (Raissi et al., 2019) without going through the discretization steps.
Our contribution includes :
@@ 2 Motivation
Let us first motivate our objective with a simple example. Physics-informed neural networks (PINNs) are a promising tool for optimizing neural networks in an unsupervised way to solve partial differential equations (PDEs). However, these methods are notoriously difficult to train (Krishnapriyan et al., 2021; Ryck et al., 2023). As an illustrative example of this challenge, let us solve the zero-boundary Poisson equation in $1$d on $\Omega=[-\pi,\pi]$.
@@
This implies that the condition number of $A$ increases extremely rapidly in the ratio between the highest and lowest frequency of the network. Given that the rate of convergence to the optimum $\Theta^{\ast}=\Theta_{0}+A^{-1}b$ can be bounded as
$\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\left(1-{c}/{\kappa(A)}\right)^{l}\|\Theta_{0}-\Theta^{\ast}\|_{2}.$ (4)
@@
We believe that this simple example clearly illustrates and highlights the fact that PINNs–even when considering a linear basis, and when the PDE is linear–suffer heavily from ill-conditioning: if $500$ steps are required in order to achieve a given error when $K=5$, roughly speaking, 312 500312500312\>500312 500 steps are required for only $K=25$.
Thus, our objective in the following will be to accelerate the convergence of such systems in this context–as well as extend them to the non-linear setting. To do so, in the following section, we will learn how to transform the optimization problem in such a way that the number of gradient descent iterations is small. The resulting method can be seen as a standalone, iterative solver as it is not only applicable to different PDEs but can handle a wide range of initial/boundary conditions and parameters.
@@ 3.1 Problem statement
For training, we assume access to a dataset of $M$ problem instances, represented by the PDE parameters $(\gamma_{i},f_{i},g_{i})_{i=1}^{M}$ and to associated target solution $(u_{i})_{i=1}^{M}$ given on a $m$ point grid $(x_{j})_{j=1}^{m}$. The solutions $(u_{i})_{i=1}^{M}$ will be used to train the neural solver. At inference, for a new PDE instance, only the PDE parameters are provided and we do not access solution points $(u_{i})_{i=1}^{M}$.
@@ 3.2 Methodology
Physics-informed neural networks consider an ansatz $u_{\Theta}$ parametrized by some finite-dimensional $\Theta$. The parameters $\Theta$ are iteratively updated by minimizing a criterion $\mathcal{L}_{\textnormal{PDE}}$ (e.g. the PDE residual), which assesses how well the ansatz $u_{\Theta}$ meets the conditions specified in equations 6 and 7. As introduced for example for PINNs (Raissi et al., 2019) or Deep Galerkin method Sirignano & Spiliopoulos (2018), we consider $\mathcal{L}_{\textnormal{PDE}}$ to be given by the strong formulation of the residual $\mathcal{L}_{\textnormal{Res}}$, plus a boundary discrepancy term $\mathcal{L}_{\textnormal{BC}}$ [^4] Note that other formulations of the loss may also be considered in a straightforward manner..
$\mathcal{L}_{\textnormal{PDE}}=\mathcal{L}_{{}_{\textnormal{Res}}}+\lambda\mathcal{L}_{\textnormal{BC}},\quad\mathcal{L}_{{}_{\textnormal{Res}}}=\sum_{x_{j}\in\Omega}|\mathcal{N}(u_{\Theta};\gamma)(x_{j})-f(x_{j})|^{2},\quad\mathcal{L}_{\textnormal{BC}}=\sum_{x_{j}\in\partial\Omega}|\mathcal{B}(u_{\Theta})(x_{j})-g(x_{j})|^{2}$ (8)
As illustrated in section 2, performing gradient descent, or alternatives such as Adam and L-BFGS on such a highly ill-conditioned loss $\mathcal{L}_{\textnormal{PDE}}$ leads to severe training difficulties (Krishnapriyan et al., 2021). The key idea in our work is to improve baseline gradient descent algorithms with the neural solver. More precisely starting from a baseline gradient algorithm, SGD in our instantiation, instead of considering the classical update, we first transform the gradient using a neural network $\mathcal{F}_{\varrho}$ with parameters $\varrho$, depending on the values of the PDE parameters as well as on other inputs such as the residual gradient provided by SGD: $\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}$. The objective is to transform, through the neural solver $\mathcal{F}_{\varrho}$, the ill-conditioned problem into a new, simpler problem that requires fewer steps in order to achieve a given error.
@@
Designed as a parametric PDE solver, $\mathcal{F}_{\varrho}$ is trained from input target data from different sets of PDE parameters, as outlined in section 3.1. Once trained, it will be used without retraining on new instances of the PDE, i.e. with new values of the PDE parameters. The underlying hypothesis is that even though the solutions may be different for different inputs and parameters, the solution methodology remains relatively consistent. This consistency is expected to enhance the algorithm’s ability to generalize across novel scenarios effectively.
@@ 3.3 Training of a physics-informed solver
Choice of Ansatz $u_{\Theta}$. A very common choice (Shen et al., 2011) is to consider a family of basis functions $\Psi(x)=\{\psi_{i}(x)\}_{i=1}^{N}$ and consider the ansatz to be given by its linear span ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$. In the following, we consider this linear reconstruction, although our formulation is generic in the sense that it can also accommodate nonlinear variants.  [^5] Although in our preliminary results we have found this may further complicate training.
As indicated in eq. 10, the solver $\mathcal{F}_{\varrho}$ will be trained from samples of the PDE parameter distribution $(\gamma,f,g)$ and from the associated samples of the solution $u$. We first describe below the inference step aiming at iteratively updating the parameters $\Theta$ of the solution function $u_{\Theta}$ while $\mathcal{F}_{\varrho}$ is hold fixed. We then describe how the ${\varrho}$ parameters of the solver are trained. Please refer to fig. 1.
@@
Algorithm 1 Inference
Inference The inference step is performed at fixed values of the solver parameters $\mathcal{F}_{\varrho}$. It consists, for a given instance of the PDE with parameters $(\gamma_{i},f_{i},g_{i})$, in finding the best $\Theta$ with a few steps of the solver using eq. 9. It is illustrated in fig. 1 - grey box: starting from initial parameters $\Theta_{0}$, we compute the Physical loss $\mathcal{L}_{\textnormal{PDE}}$ using the ansatz $u_{\Theta_{0}}$. The PDE derivatives, in $\mathcal{L}_{\textnormal{PDE}}$, can be computed by hand or automatic differentiation depending on the application  [^6] In our experiments, we computed the derivative by hand when possible since it fastens computations.. Then, eq. 9 is used to update the parameters $\Theta$ for a given number of steps $L$. The final solution is reconstructed using the linear combination, ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$, introduced in section 3.2 with the computed coefficients $\Theta_{L}$. Note that inference does not make use of the sampled target solutions $(u_{i})_{i=1}^{M}$ computed on the grid points$(x_{j})_{j=1}^{m}$.These targets are used exclusively for training the neural solver.
@@
             Compute $\Theta_{L}=\textnormal{Solve}(\Theta_{0},\gamma,f,g)$ using Algorithm 1
@@
Training the neural solver Training amounts at learning the parameters of the solver $\mathcal{F}_{\varrho}$ and is performed with a training set of PDE parameters and simulation data considered as ground truth $(\gamma_{i},f_{i},g_{i},u_{i})_{i=1}^{M}$, corresponding to PDE instances (i.e. with different parameters $\gamma$ and/or forcing terms $f$ and/or initial/boundary conditions $g$). See fig. 1 - white box. The objective is to learn a solver $\mathcal{F}_{\varrho}$ able, at inference, to converge to a target solution in a small ($2$ to $5$ in our experiments) number of steps. For that, an optimizer (Adam in our experiments) is used to update the $\mathcal{F}_{\varrho}$ parameters. The training algorithm makes use of the data associated to the different PDE instances by sampling PDEs in batches and running algorithm 1 on several PDE instances. For each PDE instance, one starts from an initial parameter value $\Theta_{0}$ and then performs two optimisation steps (see algorithm 2): (i) one consists in solving in the ansatz parameters $\Theta$ using the neural solver using algorithm 1, leading to $\Theta_{L}$; the second one is the optimization of the solver parameters $\varrho$. We train the outputs directly to match the associated ground truth $(u_{i}(x_{j}))_{j=1}^{m}$ using the data loss:
@@
The expectation is computed on the distribution of the PDE parameters $(\gamma,f,g)$. The solution $u$ is entirely determined by these parameters as indicated by the notation $u_{\gamma,f,g}$. $||u_{\Theta_{L}}-u||$ denotes a distance between the target $(u(x_{j}))_{j=1}^{m}$ and the forecast $(u_{\Theta_{L}}(x_{j}))_{j=1}^{m}$ with $m$ the trajectory size. In practice, one samples a set of PDE instances $(\gamma_{i},f_{i},g_{i})$ and for each instance a corresponding sample $u_{i}$.
Theoretical analysis and relation to preconditioning Analyzing the behavior of the inference algorithm is challenging due to the non linear nature of the solver. We however could get some intuition using simplifying assumptions. We build on the ideas introduced in section 2. We only sketch the main insights here and refer to appendix B for a proof and a more detailed analysis.
  • Using a linearization of the neural solver it can be shown that the solver performs as a pre-conditioner on the linear system.
@@
  • In practice, and as shown in the experiments, this suggests that the convergence rate is improved w.r.t. the reference baseline gradient algorithm. In the experiments section, we provide experimental comparisons with L-BFGS algorithm, which can be seen as a non linear pre-conditioning technique Rathore et al. (2024).

### v3

# Learning a Neural Solver for Parametric PDEs to Enhance Physics-Informed Methods

@@ Abstract
Physics-informed deep learning often faces optimization challenges due to the complexity of solving partial differential equations (PDEs), which involve exploring large solution spaces, require numerous iterations, and can lead to unstable training. These challenges arise particularly from the ill-conditioning of the optimization problem caused by the differential terms in the loss function. To address these issues, we propose learning a solver, i.e., solving PDEs using a physics-informed iterative algorithm trained on data. Our method learns to condition a gradient descent algorithm that automatically adapts to each PDE instance, significantly accelerating and stabilizing the optimization process and enabling faster convergence of physics-aware models. Furthermore, while traditional physics-informed methods solve for a single PDE instance, our approach extends to parametric PDEs. Specifically, we integrate the physical loss gradient with PDE parameters, allowing our method to solve over a distribution of PDE parameters, including coefficients, initial conditions, and boundary conditions. We demonstrate the effectiveness of our approach through empirical experiments on multiple datasets, comparing both training and test-time optimization performance. The code is available at https://github.com/2ailesB/neural-parametric-solver.
@@ 1 Introduction
Partial Differential Equations (PDEs) are ubiquitous as mathematical models of dynamical phenomena in science and engineering. Solving PDEs is of crucial interest to researchers and engineers, leading to a huge literature on this subject (Evans, 2010; Salsa, 2015). Traditional approaches to solving PDEs such as finite difference, finite element analysis, or spectral methods (Zienkiewicz et al., 2005; LeVeque, 2007) often come with stability and convergence guarantees but suffer from a high computational cost. Improving numerical PDE solvers through faster and more accurate algorithms remains an active research topic (Zienkiewicz et al., 2005).
@@
PDE solvers usually rely on discretization and/or linearization of the problem through various techniques to simplify the computations. Iterative methods such as Jacobi, Gauss-Seidel, Conjugate Gradient, and Krylov subspace methods can then be used to solve the resulting systems. Unfortunately, many PDEs have an ill-conditioned nature, and these iterative processes can demand extensive computational resources. Preconditioning techniques are often essential to mitigate this, though they require precise customization to the specific PDE problem, making the development of effective solvers a significant research endeavor in itself. Yet, the computational demands, time, and expertise required to develop these algorithms sometimes make them impractical or sub-optimal for specific classes of problems. Instead of relying on hand-designed algorithms, researchers have investigated, as an alternative, the use of machine learning to train iterative PDE solvers Hsieh et al. (2019); Li et al. (2023a); Rudikov et al. (2024); Kopaničáková et al. (2023). These approaches usually parallel the classical numerical methods by solving a linear system resulting from the discretization of a PDE, for example, using finite differences or finite elements. A preconditioner is learned from data by optimizing a residual loss computed w.r.t. a ground truth solution obtained with a PDE solver. This preconditioner is used on top of a baseline iterative solver and aims at accelerating its convergence. Examples of baseline solvers are the conjugate gradient (Li et al., 2023a; Rudikov et al., 2024) or the Jacobi method Hsieh et al. (2019).
@@
Another recent research direction investigates the use of neural networks for building surrogate models in order to accelerate the computations traditionally handled by numerical techniques. These methods fall into two main categories: supervised and unsupervised. The supervised methodology consists of first solving the PDE using numerical methods to generate input and target data and then regressing the solution using neural networks in the hope that this surrogate could solve new instances of the PDE. Many models, such as Neural Operators, lie within this class (Li et al., 2020; Raonić et al., 2023; Bartolucci et al., 2023) and focus on learning the solution operator directly through a single neural network pass. Unsupervised approaches, involve considering a neural network as a solution of the PDE. The neural network parameters are found by minimizing the PDE residual with gradient descent. Methods such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), or DeepRitz (E & Yu, 2018) fall under this category. This family of methods is attractive as it does not rely on any form of data, but only on information from the PDE residual. However, they exhibit severe difficulties during training (Krishnapriyan et al., 2021; Ryck et al., 2023), often requiring many optimization steps and sophisticated training schemes (Krishnapriyan et al., 2021; Rathore et al., 2024). The ill-conditioned nature of PDE residual loss appears again in this context, making standard optimizers such as Adam inappropriate (see section E.6 and section E.7 for a visualization of the ill conditioning of this loss landscape). A detailed review of the existing literature is described in appendix A.
@@
In this work, we consider having access to the PDE as in unsupervised approaches and also to some data for training our neural solver. Our objective is to solve the optimization issues mentioned above by learning an iterative algorithm that solves the PDE from its residual, defined as in the PINNs framework (see fig. 1). This neural solver is trained from data, either simulations or observations. Different from the classical ML training problem which aims at learning the parameters of a statistical model from examples, the problem we handle is learning to learn, Andrychowicz et al. (2016), i.e. learning an iterative algorithm that will allow us to solve a learning problem. When vanilla PINNs handle a single PDE instance, requiring retraining for each new instance, we consider the more complex setting of solving parametric PDEs, the parameters may include boundary or initial conditions, forcing terms, and PDE coefficients. Each specific instance of the PDE, sampled from the PDE parameter distribution, will then be considered as a training example. The objective is then to learn a solver from a sample of the parametric PDE distribution in order to accelerate inference on new instances of the equation. With respect to unsupervised approaches, our model implements an improved optimization algorithm, tailored to the parametric PDE problem at hand instead of using a hand-defined method such as stochastic gradient descent (SGD) or Adam. As demonstrated in the experimental section, this approach proves highly effective for the ill-posed problem of optimizing the PINNs objective, enabling convergence in just a few steps. This is further illustrated through gradient trajectory visualizations in section E.6. In the proposed methodology, the neural solver will make use of the gradient information computed by a baseline gradient method to accelerate its convergence. In our instantiation, we will use SGD as our baseline algorithm, but the method could be easily extended to other baselines. Our model deviates from the traditional preconditioning methods by directly optimizing the non-linear PDE residual loss (Raissi et al., 2019) without going through the discretization steps. Our contribution includes :
@@ 2 Motivation
Let us first motivate our objective with a simple example. Physics-informed neural networks (PINNs) are a promising tool for optimizing neural networks in an unsupervised way to solve partial differential equations (PDEs). However, these methods are notoriously difficult to train (Krishnapriyan et al., 2021; Ryck et al., 2023). As an illustrative example of this challenge, let us solve the zero-boundary Poisson equation in $1$d on $\Omega=[-\pi,\pi]$. Note that this section is intentionally informal, we provide rigorous statements and proofs in Appendix B.
@@
This implies that the condition number of $A$ increases extremely rapidly in the ratio between the highest and lowest frequencies of the network. Given that the rate of convergence to the optimum $\Theta^{\ast}=\Theta_{0}+A^{-1}b$ can be bounded as
$\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\left(1-{c}/{\kappa(A)}\right)^{l}\|\Theta_{0}-\Theta^{\ast}\|_{2},$ (4)
@@
We believe that this simple example clearly illustrates and highlights the fact that PINNs–even when considering a linear basis, and when the PDE is linear–suffer heavily from ill-conditioning: if $500$ steps are required in order to achieve a given error when $K=5$, roughly speaking, 312 500312500312\>500312 500 steps are required for only $K=25$. This result extends to more general linear systems of equations and linear ansatz, as explained in appendix B.
Our objective in the following will be to accelerate the convergence of such systems in this context–as well as extend them to the non-linear setting. To do so, in the following section, we will learn how to transform the optimization problem in such a way that the number of gradient descent iterations is small. The resulting method can be seen as a standalone, iterative solver as it is not only applicable to different PDEs but can handle a wide range of initial/boundary conditions and parameters.
@@ 3.1 Problem statement
For training, we assume access to a dataset of $M$ problem instances, represented by the PDE parameters $(\gamma_{i},f_{i},g_{i})_{i=1}^{M}$ and to associated target solution $(u_{i})_{i=1}^{M}$ given on a $m$ point grid $(x_{j})_{j=1}^{m}$. The solutions $(u_{i})_{i=1}^{M}$ will be used to train the neural solver. At inference, for a new PDE instance, only the PDE parameters are provided and we do not have access to solution points $(u_{i})_{i=1}^{M}$.
@@ 3.2 Methodology
Physics-informed neural networks consider an ansatz $u_{\Theta}$ parametrized by some finite-dimensional $\Theta$. The parameters $\Theta$ are iteratively updated by minimizing a criterion $\mathcal{L}_{\textnormal{PDE}}$ (e.g. the PDE residual), which assesses how well the ansatz $u_{\Theta}$ meets the conditions specified in equations 6 and 7. As introduced for example for PINNs (Raissi et al., 2019) or Deep Galerkin method (Sirignano & Spiliopoulos, 2018), we consider $\mathcal{L}_{\textnormal{PDE}}$ to be given by the strong formulation of the residual $\mathcal{L}_{\textnormal{Res}}$, plus a boundary discrepancy term [^3] Note that other formulations of the loss may also be considered in a straightforward manner. $\mathcal{L}_{\textnormal{BC}}$: $\mathcal{L}_{\textnormal{PDE}}=\mathcal{L}_{{}_{\textnormal{Res}}}+\lambda\mathcal{L}_{\textnormal{BC}},\quad,\lambda>0$.
$\mathcal{L}_{{}_{\textnormal{Res}}}=\sum_{x_{j}\in\Omega}|\mathcal{N}(u_{\Theta};\gamma)(x_{j})-f(x_{j})|^{2},\quad\mathcal{L}_{\textnormal{BC}}=\sum_{x_{j}\in\partial\Omega}|\mathcal{B}(u_{\Theta})(x_{j})-g(x_{j})|^{2}$ (8)
As illustrated in section 2, performing gradient descent, or alternatives such as Adam and L-BFGS on such a highly ill-conditioned loss $\mathcal{L}_{\textnormal{PDE}}$ leads to severe training difficulties (Krishnapriyan et al., 2021). The key idea in our work is to improve a baseline gradient descent algorithms with the neural solver. More precisely starting from a baseline gradient algorithm, SGD in our instantiation, instead of considering the classical update, we first transform the gradient using a neural network $\mathcal{F}_{\varrho}$ with parameters $\varrho$, depending on the values of the PDE parameters as well as on other inputs such as the residual gradient provided by SGD: $\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}$. The objective is to transform, through the neural solver $\mathcal{F}_{\varrho}$, the ill-conditioned problem into a new, simpler problem that requires fewer steps in order to achieve a given error.
@@
Designed as a parametric PDE solver, $\mathcal{F}_{\varrho}$  [^4] In the following we use ”solver” with $\mathcal{F}_{\varrho}$ as a short-hand to refer to our proposed method. is trained with input target data from different sets of PDE parameters, as outlined in section 3.1. Once trained, it will be used without retraining on new instances of the PDE, i.e. with new values of the PDE parameters. The underlying hypothesis is that even though the solutions may be different for different inputs and parameters, the solution methodology remains relatively consistent. This consistency is expected to enhance the algorithm’s ability to generalize across novel scenarios effectively.
@@ 3.3 Training of a physics-informed solver
Choice of Ansatz $u_{\Theta}$. A very common choice (Shen et al., 2011) is to consider a family of basis functions $\Psi(x)=\{\psi_{i}(x)\}_{i=1}^{N}$ and consider the ansatz to be given by its linear span ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$. In the following, we consider this linear reconstruction, although our formulation is generic in the sense that it can also accommodate nonlinear variants.  [^5] Although we have found this may further complicate training.
As indicated in eq. 10, the solver $\mathcal{F}_{\varrho}$ will be trained from samples of the PDE parameter distribution $(\gamma,f,g)$ and from the associated samples of the solution $u$. We first describe below the inference step aiming at iteratively updating the parameters $\Theta$ of the solution function $u_{\Theta}$ while $\mathcal{F}_{\varrho}$ is held fixed. We then describe how the ${\varrho}$ parameters of the solver are trained. Please refer to fig. 1 that illustrates the interaction between the two steps.
@@
Algorithm 1 Inference using the neural PDE solver.
@@
             Estimate $\Theta_{L}$ from $\Theta_{0},(\gamma,f,g)$ using algorithm 1
@@
             Update $\varrho$ with gradient descent from the data loss in eq. 10
@@
Inference The inference step is performed at fixed values of the solver parameters $\mathcal{F}_{\varrho}$. It consists, for a given instance of the PDE with parameters $(\gamma_{i},f_{i},g_{i})$, in finding the best $\Theta$ with a few steps of the solver using eq. 9. It is illustrated in fig. 1 - grey box and formalized in algorithm 1: starting from initial parameters $\Theta_{0}$, we compute the Physical loss $\mathcal{L}_{\textnormal{PDE}}$ using the ansatz $u_{\Theta_{0}}$. The PDE derivatives in $\mathcal{L}_{\textnormal{PDE}}$, can be computed by hand or automatic differentiation depending on the application  [^6] In our experiments, we computed the derivative by hand when possible since it fastens computations.. Then, eq. 9 is used to update the parameters $\Theta$ for a given number of steps $L$. The final solution is reconstructed using the linear combination, ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$, introduced in section 3.2 with the computed coefficients $\Theta_{L}$. Note that inference does not make use of the sampled target solutions $(u_{i})_{i=1}^{M}$ computed on the grid points$(x_{j})_{j=1}^{m}$.These targets are used exclusively for training the neural solver.
Training the neural solver Training amounts to learning the parameters of the solver $\mathcal{F}_{\varrho}$ and is performed with a training set of PDE parameters and simulation data considered as ground truth $(\gamma_{i},f_{i},g_{i},u_{i})_{i=1}^{M}$, corresponding to PDE instances (i.e. with different parameters $\gamma$ and/or forcing terms $f$ and/or initial/boundary conditions $g$). See fig. 1 - white box and algorithm 2. The objective is to learn a solver $\mathcal{F}_{\varrho}$ able, at inference, to converge to a target solution in a small ($2$ to $5$ in our experiments) number of steps. For that, an optimizer (Adam in our experiments) is used to update the $\mathcal{F}_{\varrho}$ parameters. The training algorithm makes use of the data associated with the different PDE instances by sampling PDEs in batches and running algorithm 1 on several PDE instances. For each PDE instance, one starts from an initial parameter value $\Theta_{0}$ and then performs two optimization steps (see algorithm 2): (i) one consists in solving in the ansatz parameters $\Theta$ using the neural solver using algorithm 1, leading to $\Theta_{L}$; (ii) the second one is the optimization of the solver parameters $\varrho$. We train the outputs directly to match the associated ground truth $(u_{i}(x_{j}))_{j=1}^{m}$ using the data loss:
@@
The expectation is computed on the distribution of the PDE parameters $(\gamma,f,g)$. The solution $u$ is entirely determined by these parameters as indicated by the notation $u_{\gamma,f,g}$. $||u_{\Theta_{L}}-u||$ denotes a distance between the target $(u(x_{j}))_{j=1}^{m}$ and the forecast $(u_{\Theta_{L}}(x_{j}))_{j=1}^{m}$ with $m$ the trajectory size  [^7] To simplify the notation, we used a fixed grid size $m$. However, this framework can be used with different grid sizes, as well as irregular grids. See Ablation in appendix E, table 13. In practice, one samples a set of PDE instances $(\gamma_{i},f_{i},g_{i})$ and for each instance a corresponding sample $u_{i}$.
Theoretical analysis and relation to preconditioning
Analyzing the behavior of the inference algorithm is challenging due to the non linear nature of the solver. We however could get some intuition using simplifying assumptions. We build on the ideas introduced in section 2 for the simple case of the Poisson equation, for which an explicit analytical solution could be derived. We provide in appendix B, a proof for a more general case and give below in theorem 1 our main result. This shows that the number of steps induced by $\mathcal{F}$ for the proposed algorithm is significantly less than the number of steps required by the baseline PINNs algorithm. This results is obtained under two main assumptions: (i) $\mathcal{F}$ behaves like its linearization and (ii) the descent operator $\mathcal{F}$ used in our algorithm, allows us to reach the optimum of $\mathcal{L}_{\text{DATA}}$.
Theorem 1.
(Convergence rate in the linear case). Given a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, assume the conditioner $\mathcal{F}$ behaves like its linearization $P=\text{Jacobian}(\mathcal{F})$, meaning that $\mathcal{F}$ can be replaced by $P$ at any point. Let A be the matrix derived from the PDE loss as eq. 3 for the Poisson equation or eq. 15 in the more general case. Denote by $\kappa(A)$ the condition number of the matrix $A$. The number of steps $N^{\prime}(\varepsilon)$ required to achieve an error $\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\varepsilon$ satisfies:
$N^{\prime}(\varepsilon)=O\left(\kappa(PA)\ln\left(\tfrac{1}{\varepsilon}\right)\right),$ (11)
Moreover, if $\mathcal{F}$ minimizes $\mathcal{L}_{\text{DATA}}$ this necessarily implies $\kappa(PA)=1\leq\kappa(A)$. Consequently, the number of steps is effectively reduced, i.e., $N^{\prime}(\varepsilon)\ll N(\varepsilon)$ with $N(\varepsilon)$ the number of steps of the vanilla PINNs.
Proof.
We sketch the main insights here and refer to appendix B for the proof and a detailed analysis.
  • Using a linearization of the neural solver, it can be shown that the solver performs as a pre-conditioner on the linear system.
@@
  • In practice, and as shown in the experiments (section 4), the convergence rate is significantly improved w.r.t. the reference baseline gradient algorithm.